OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______ Please explain!

OpenStudy (anonymous):

f(5) is 10 because 2*1 = 2

OpenStudy (freckles):

Try to find f(4)

OpenStudy (freckles):

f(3)=f(1)+f(2)+f(3-1)=f(1)+f(2)+f(2)=2+5+5=12

OpenStudy (anonymous):

what?

OpenStudy (freckles):

You have to find f(3) then f(4) then f(5)

OpenStudy (freckles):

I found f(3) for you and I want you to try to find f(4)

OpenStudy (freckles):

Remember you have f(n)=f(1)+f(2)+f(n-1)

OpenStudy (freckles):

replace the n's with 3's you have f(3)=f(1)+f(2)+f(3-1)

OpenStudy (freckles):

This means we can find f(3) since we know the values for f(1) and f(2)

OpenStudy (freckles):

f(3)=f(1)+f(2)+f(2) =2 +5 +5 I replaced f(2) with 5 because your problem says f(2) is 5 I replaced f(1) with 2 because your problem says f(1) is 2

OpenStudy (anonymous):

plz explain

OpenStudy (freckles):

Now you can find f(4) replace all the n's again in f(n)=f(1)+f(2)+f(n-1) we are going to replace n with 4 this time since we want f(4) f(n)=f(1)+f(2)+f(n-1) replace n with 4 f(4)=f(1)+f(2)+f(4-1) f(4)=f(1)+f(2)+f(3) f(4)=2 + 5 + 12 f(4)=7+12 f(4)=19 I will leave the last step for you which is finding f(5)

OpenStudy (freckles):

Have you tried to find f(5) yet?

OpenStudy (anonymous):

I still don't understand. :(

OpenStudy (anonymous):

how did u know n was 4

OpenStudy (freckles):

n varies n can equal 3,4,5,6,7,8,9,10, and so on...

OpenStudy (freckles):

f(1) is the first number in the sequence f(2) is the second number in the sequence f(3) is the third number in the sequence f(4) is the fourth number in the sequence f(5) is the fifth number in the sequence

OpenStudy (freckles):

f(1) and f(2) are given and f(n)=f(1)+f(2)+f(n-1) equals for every integer n greater than or equal to 3

OpenStudy (freckles):

To find f(3) we need to replace n with 3 so we have f(3)=f(1)+f(2)+f(3-1) =f(1)+f(2)+f(2) =2 +5 +5 =12 So now that you have f(3) you can find f(4) by just replacing the n with 4 now f(4)=f(1)+f(2)+f(4-1) =f(1)+f(2)+f(3) = 2 + 5 +12 =19 Now you can find f(5) just by replacing the n with 5 f(5)=f(1)+f(2)+f(5-1) you can finish this

OpenStudy (anonymous):

f(3)=f(1)+f(2)+f(3-1) =f(1)+f(2)+f(2) =2 +5 +5 were did the five come from? =12